# Why These Documents? Explainable Generative Retrieval with Hierarchical Category Paths

### v1

@@ 4.2. Optimization with Path-Augmented Dataset
The second step is to augment the existing training set $\mathcal{X}$ with path information, building a path-augmented training set $\mathcal{X}^{+}=\{(q,p^{q},D,d)\}$. To this end, we first (1) link each query to one of the document’s candidate paths based on semantic similarity computed by pre-trained text encoder, and then (2) use the query-path pairs along with the document-path pairs to optimize the generative retrieval model.
4.2.1. Linking Path with Query.
@@
Using the candidate path set for each document, we build a path augmented training set $\mathcal{X}^{+}$. For each query-document pair in the training set $(q,D,d)\in\mathcal{X}$, we link the query $q$ to its most relevant path among the paths in the document’s candidate path set ${P}_{D}$. This query-path linking can be described as follows:
(6) $p^{q}=\operatorname*{argmax}_{p\in\mathcal{P}_{D}}\text{ sim}(E(q),E(p)),$
@@
4.2.2. Optimization.
By leveraging the path-augmented training set $\mathcal{X}^{+}$, we train our model $\mathcal{M}^{\theta}$ on both indexing and retrieval tasks, as described in 3.1. That is, the model is optimized by indexing loss and retrieval loss. First, the indexing loss is defined by
(7) $\mathcal{L}_{\text{indexing}}=-\sum_{p\in\mathcal{P}_{D}}\log\mathcal{M}^{\theta}(p,d\mid D)$
where $p$ is a path from $D$’s candidate path set. Note that there are multiple paths assigned to the document, so we optimize our model by using all candidate paths in $\mathcal{P}_{D}$. Through this process, the model can effectively memorize the content of the documents while mapping them to their relevant category paths. Next, the retrieval loss is described as follows:
(8) $\mathcal{L}_{\text{retrieval}}=-\log\mathcal{M}^{\theta}(p,d\mid q)$
The retrieval loss optimizes the model to learn the relationship between the query and its relevant path, allowing it to retrieve the documents that match the semantic of the query with the help of assigned paths. Finally, we define our total loss as the sum of the indexing loss and the retrieval loss, regarding the training process as a multi-task learning framework:
(9) $\mathcal{L}=\mathcal{L}_{\text{indexing}}+\mathcal{L}_{\text{retrieval}}.$
4.3. Inference with Path-Aware Reranking
@@
(10) $Y_{j}=\{(d_{i},s_{i})\sim\ \mathcal{M}^{\theta}(\cdot\mid q,p_{j})\}_{i=1}^{m},$
where $s_{i}$ represents the score for the docid $d_{i}$ conditioned on the category path $p_{j}$. The remaining process is to aggregate $K_{p}$ number of docid-score pair sets for making the final ranked list of docids. At this point, we employ path-aware reranking strategy to remain only unique docid with the highest score, resulting in $\tilde{Y}$.
(11) $\tilde{Y}=\left\{(d,s)\mid s==\max\{s^{\prime}|(d,s^{\prime})\in Y_{j}\},\forall(d,s)\in\cup_{j=1}^{K_{p}}Y_{j}\right\}$
From the set of unique docid-score pairs, we obtain the final ranked list by sorting their scores in descending order, $Y_{\text{final}}=\text{sort}(\tilde{Y})$. By utilizing path-aware reranking strategy, HyPE can effectively capture the semantic information of an input query from multiple category paths, leading to improved retrieval performance.
5. Experiments
@@
  • RQ1: Can HyPE effectively explain the semantic relationship between the query and document for document retrieval tasks?
  • RQ2: Can HyPE provide more effective and reasonable explanations compared to other methods for the user’s query?
  • RQ3: Can HyPE enhance the retrieval accuracy?
  • RQ4: Can path-aware reranking strategy of HyPE improve retrieval accuracy?
Table 1. Semantic relevance between query/explanation and document/explanation on 1,000 query-document pairs sampled from the NQ320K dev set, where each baseline successfully retrieves the ground truth document at rank 1. We also report BM25 as a sparse retrieval baseline.
@@
Figure 3. Examples of the document annotated for multiple queries in the NQ320K dev set. The generative retrieval model with HyPE generates query-specific category paths based on the topics of the document associated with each query, explaining why the document is retrieved for the particular query.
5.1. Experimental Settings
5.1.1. Dataset
We conduct our experiments on two public datasets, NQ320K (Kwiatkowski et al., 2019) and MS MARCO (Nguyen et al., 2016), which have been widely utilized in previous works (Tay et al., 2022; Wang et al., 2022). NQ320K is a dataset derived from Google’s Natural Questions (NQ) and comprises 307,373 query-document pairs along with 7,830 test queries, all linked to 109,739 documents sourced from Wikipedia. We divide the test set into two subsets, seen and unseen, following the setup in  (Wang et al., 2022; Sun et al., 2023), where the seen test includes queries whose annotated target documents are present in the training set, and the unseen test consists of queries with no labeled documents in the training set. On the other hand, MS MARCO comprises queries and web pages sourced from Bing search. We use the MS MARCO document ranking dataset, which contains 366,235 query-document relevant pairs and 5,187 test queries with 323,569 documents. The detailed statistics of the datasets are provided in Appendix 5.
5.1.2. Evaluation Metrics
We report Recall and Mean Reciprocal Rank (MRR) for NQ320K and MS MARCO. For NQ320K, we use Recall@{1, 10, 100} and MRR@100. For MS MARCO, we use Recall@{1, 10, 100} and MRR@10 as done in previous works (Sun et al., 2023; Wang et al., 2023).
Figure 4. Human evaluation of pairwise quality comparisons for retrieval explanations, generated by HyPE and baseline models (* : $p$-value ¡ 0.05).
Figure 5. An example from the NQ320K dev set showing that using only the title as the docid fails to retrieve the correct document. However, when HyPE is adopted, the generative retrieval model better understands the query and successfully retrieves the correct document.
Table 2. Retrieval accuracy of existing generative retrieval methods and our HyPE framework on the NQ320K dataset. * denotes the statistical significance on paired t-test $p<0.05$.
Method Full test Seen test Unseen test
R@1 R@10 R@100 M@100 R@1 R@10 R@100 M@100 R@1 R@10 R@100 M@100
GENRE (Cao et al., 2021) 55.2 67.3 75.4 59.9 69.5 83.7 90.4 75.0 6.0 10.4 23.4 7.8
[…]
+ HyPE 63.6 83.5 90.1 71.0 66.4 86.3 92.6 73.9 53.7 73.6 81.7 61.0
NOVO (Wang et al., 2023) 69.3 89.7 95.9 76.7 - - - - - - - -
[…]
+ HyPE 60.7 79.1 86.2 67.6 66.6 84.6 90.7 73.4 40.1 60.2 70.6 47.5
SEAL (Bevilacqua et al., 2022) 59.9 81.2 90.9 67.7 - - - - - - - -
[…]
+ HyPE 61.5 79.6 85.2 68.3 66.3 84.6 89.8 73.2 44.8 62.2 69.4 51.3
GENRET (Sun et al., 2023) 68.1 88.8 95.2 75.9 70.2 90.3 96.0 77.7 62.5 83.6 92.5 70.4
[…]
+ HyPE 64.5 84.2 90.2 71.9 69.5 88.6 93.8 76.8 47.2 68.7 77.6 55.0
5.1.3. Baselines
To validate the effectiveness of HyPE in generative retrieval, we conduct experiments on four representative docid types, introduced in Section 3.2.2, as our baseline generative retrieval approach.
  • Title docid (Cao et al., 2021) uses a document’s title as docid. For instances without a title, we use the first 16 tokens of the document as a pseudo-title, following the approach used in (Sun et al., 2023).
  • Keword docid (Wang et al., 2023) uses a sequence of keywords as docid that effectively represent the document. [^8] We use the off-the-shelf keyword extraction tool (Grootendorst, 2020). For NQ320K, we use 3 keywords, while for MS MARCO, we extract 5 keywords.
  • Summary docid uses the document summary as docid. [^9] We use the off-the-shelf text summarization model (Lewis et al., 2020). Although it has not been attempted before, a similar structure using substrings is employed in (Bevilacqua et al., 2022).
  • Atomic docid (Sun et al., 2023) uses a unique arbitrary integer as docid. We assign each document a unique arbitrary integer and generates a corresponding new token for it.
Existing generative methods employ various architectures and optimization techniques, which may introduce additional factors affecting performance. To isolate and specifically assess the impact of HyPE, we use the basic form of generative retrieval introduced in Section 3 as our main baseline. This approach allows us to focus on a direct comparison between plain docids and those enhanced with HyPE, ensuring that the observed effects are due to HyPE itself rather than other architectural or optimization variations. Furthermore, we additionally report the performance of existing methods to provide a reference for the performance of existing generative methods. We set a reference baseline for each docid type. Specifically, for each docid type, method that achieves the highest performance using the same dataset as ours is selected as the reference baseline for each docid. The detailed criteria for selecting the reference baseline can be found in 5.
5.1.4. Implementation Details
We use T5-base (Raffel et al., 2020) as our backbone model. For the input of the indexing task, we utilize the FirstP approach as our document representations (Section 3.2.1). Additionally, for the indexing task, we employ five synthetic queries, generated by using docT5query (Nogueira and Lin, 2020) with nucleus sampling with parameters $p=0.8$ and $t=0.8$. We use new [DOC] token to separate the path from the docid, which we insert between the path and the docid. We optimize our model as described in 4.2.2, while employing AdamW optimizer with a learning rate of 5e-4 and a batch size of 128, for up to 1M training steps. During the inference of HyPE, we adopt path-aware reranking strategy; for the path generation stage, we generates three category paths (i.e., $K_{p}=3$), and for the docid generation stage, we use constrained beam search with a beam size of 100 (i.e., $m=100$).
5.2. Explainability of HyPE (RQ1 & RQ2)
5.2.1. Quantitative analysis (RQ1).
We first quantitatively evaluate whether HyPE’s category path provides a valid explanation by effectively capturing the relationship between the query and the document. To do this, we use a model for the task of semantic textual similarity (STS), which measures the semantic relevance between two sentences (Agirre et al., 2012). Using the task, we assess the semantic relevance between the query and the explanation, as well as between the document and the explanation. Specifically, for each baseline, we use the model output as an explanation and calculate the STS scores for both the query-explanation and document-explanation pairs. We then compute the geometric mean of these two scores to evaluate how effectively the explanation captures the relationship between the query and the document. In this experiment, we also include BM25 as a baseline that can provide explanations for its retrieval results. For the explanation of BM25, we consider the top-3 terms with the highest BM25 scores between a query and a document.

### v2

@@ 3.2 Optimization with category path
The second step is to augment the training set $\mathcal{X}$ with path, building a path-augmented training set $\mathcal{X}^{+}=\{(q,p^{q},D,d)\}$. To achieve this, we first (1) link each query to one of the document’s candidate paths based on semantic similarity computed by pre-trained encoder, and then (2) utilize the resulting query-path pairs together with the document-path pairs to optimize the retrieval model.
Linking Path with Query.
@@
Using the candidate path set for each document, we build a path augmented training set $\mathcal{X}^{+}$. For each query-document pair in the training set $(q,D,d)\in\mathcal{X}$, we link the query $q$ to its most relevant path among the paths in the document’s candidate path set ${P}_{D}$. This linking can be described as follows:
$p^{q}=\operatorname*{argmax}_{p\in\mathcal{P}_{D}}\text{ sim}(E(q),E(p)),$ (5)
@@
Optimization.
By leveraging the path-augmented training set $\mathcal{X}^{+}$, we train our model $\mathcal{M}^{\theta}$ on both indexing and retrieval tasks, as described in 2.1. Our optimization follows the same strategy as standard generative retrieval in 2.1, with the only difference being the addition of path information as follows:
$\max_{\theta}\sum\mathcal{M}^{\theta}(p^{q},d\mid D)+\mathcal{M}^{\theta}(p^{q},d\mid q)$ (6)
3.3 Inference with Path-Aware Ranking
@@
$Y_{j}=\{(d_{i},s_{i})\sim\ \mathcal{M}^{\theta}(\cdot\mid q,p_{j})\}_{i=1}^{m},$ (7)
where $s_{i}$ represents the score for the docid $d_{i}$ conditioned on the category path $p_{j}$. The remaining process is to aggregate $K_{p}$ number of docid-score pair sets for making the final ranked list of docids. At this point, we remain only unique docid with the highest score, resulting in $\tilde{Y}$.
$\tilde{Y}=\left\{(d,s)\mid s=\max\{s^{\prime}|(d,s^{\prime})\in Y_{j}\},\forall(d,s)\in\cup_{j=1}^{K_{p}}Y_{j}\right\}$ (8)
From the set of unique docid-score pairs, we obtain the final ranked list by sorting their scores in descending order, $Y_{\text{final}}=\text{sort}(\tilde{Y})$. By utilizing path-aware ranking strategy, HyPE can effectively capture the semantic information of an input query from multiple category paths, leading to improved retrieval performance.
4 Experiments
@@
  • RQ1: Can HyPE improve retrieval accuracy?
  • RQ2: Can hierarchical category paths in HyPE serve as effective explanations for retrieval?
  • RQ3: Can explanations of HyPE help real-world users in search systems?
Method Full test Seen test Unseen test
R@1 R@10 R@100 M@100 R@1 R@10 R@100 M@100 R@1 R@10 R@100 M@100
Title docid 62.2 78.7 89.3 68.6 64.8 81.5 90.1 71.2 53.1 68.9 80.4 59.3
+ HyPE 63.6 83.5 90.1 71.0 66.4 86.3 92.6 73.9 53.7 73.6 81.7 61.0
Keyword docid 61.8 77.1 85.5 67.6 67.3 82.3 89.9 73.0 43.0 59.0 70.4 48.8
+ HyPE 60.7 79.1 86.2 67.6 66.6 84.6 90.7 73.4 40.1 60.2 70.6 47.5
Summary docid 60.9 78.8 84.1 67.6 65.7 84.1 88.6 72.6 44.0 60.5 68.5 50.1
+ HyPE 61.5 79.6 85.2 68.3 66.3 84.6 89.8 73.2 44.8 62.2 69.4 51.3
Atomic docid 65.3 83.5 89.3 72.2 70.2 88.3 93.5 77.2 48.6 66.8 74.9 55.0
+ HyPE 64.5 84.2 90.2 71.9 69.5 88.6 93.8 76.8 47.2 68.7 77.6 55.0
Table 1: Retrieval accuracy of baselines and our HyPE framework on the NQ320K. * denotes the statistical significance on paired t-test $p<0.05$.
Method R@1 R@10 R@100 M@10
Keyword docid 31.7 61.2 77.2 41.0
+ HyPE 32.2 62.7 78.5 41.9
Improvement +1.6% +2.5% +1.7% +2.2%
Summary docid 28.1 55.5 71.5 36.8
+ HyPE 28.4 57.5 73.1 37.8
Improvement +1.1% +3.6% +2.2% +2.7%
Atomic docid 43.9 73.6 85.6 53.8
+ HyPE 44.9 74.6 87.1 54.7
Improvement +2.3% +1.4% +1.8% +1.7%
Table 2: Retrieval accuracy of baselines and HyPE on the MS MARCO. * denotes the statistical significance on paired t-test $p<0.05$.
4.1 Experimental Settings
Dataset.
We conduct our experiments on two datasets, NQ320K Kwiatkowski et al. (2019) and MS MARCO Nguyen et al. (2016), which have been widely utilized in previous works Tay et al. (2022); Wang et al. (2022). For NQ320K, we divide the test set into two subsets, seen and unseen, following the setup in  Wang et al. (2022); Sun et al. (2023), where the seen test includes queries whose annotated target documents are present in the training set, and the unseen test consists of queries with no labeled documents in the training set. More details are provided in Appendix A.4.
Evaluation Metrics.
We report Recall and Mean Reciprocal Rank (MRR) for NQ320K and MS MARCO. For NQ320K, we use Recall@{1, 10, 100} and MRR@100. For MS MARCO, we use Recall@{1, 10, 100} and MRR@10 as done in previous works Sun et al. (2023); Wang et al. (2023).
Baselines.
To validate the effectiveness of HyPE across diverse generative retrieval settings, we conduct experiments on four representative docid types, introduced in Section 2.2, as our baseline.
  • Title docid uses a document’s title as docid. For documents without a title, we use the first 16 tokens of the document as a title, following the approach used in Sun et al. (2023).
  • Keyword docid uses a sequence of keywords as docid that effectively represent the document. For NQ320K, we use 3 keywords, while for MS MARCO, we extract 5 keywords.
  • Summary docid uses the document summary as docid. Although it has not been attempted before, a similar structure using substrings is employed in Bevilacqua et al. (2022).
  • Atomic docid uses a unique arbitrary integer as docid. We assign each document a integer and generates a corresponding new token for it.
We intentionally do not consider semantic docids (+HyPE) in our experiments. This is because semantic docids are constructed based on techniques such as hierarchical clustering, and thus inherently embed a semantic structure. Given that these structures are already formed in a coarse-to-fine manner, prepending hierarchical category paths to them can contradict the coarse-to-fine principle.
Furthermore, existing generative methods employ various architectures and optimization techniques, which may introduce additional factors affecting performance. To specifically assess the impact of HyPE, we adopt the basic form of generative retrieval described in Section 2 as our baseline. This approach ensures a direct comparison between plain docids and those enhanced with HyPE, isolating the effects of HyPE itself from other architectural or optimization differences. For more details, please refer to the Appendix A.7.
4.1.1 Implementation Details
We use T5-base Raffel et al. (2020) as our backbone model. For the input of the indexing task, we utilize the FirstP approach as our document representations and five synthetic queries. (Section 2.2). During the inference of HyPE, we generate three category paths (i.e., $K_{p}=3$), and for the docid decoding stage, we use constrained beam search with a beam size of 100 (i.e., $m=100$). More details about this part are provided in Appendix A.7.
4.2 HyPE improves retrieval accuracy (RQ1)
Table 1 shows retrieval accuracy of various docid types with HyPE on NQ320K. Overall, HyPE consistently improves retrieval accuracy across all docid types in both seen test and unseen test. This demonstrates that HyPE’s hierarchical category paths can be orthogonally applied to enhance retrieval accuracy across different docid types, suggesting that integrating these paths into existing generative retrieval methods can further improve performance. While HyPE can be applied to all docid types effectively, the experimental results show that title docid yields the most significant performance improvement when HyPE is applied. Our paths, serve as a pseudo-reasoning, allowing the model to navigate step-by-step through various semantic hierarchical categories before arriving docid. Since titles are concise and inherently reflect a structured overview of a document, they aligns well with the HyPE’s hierarchical category paths, further enhancing retrieval accuracy.
Figure 3: Human evaluation of pairwise quality comparisons for retrieval explanations, generated by HyPE and baseline models.
Additionally, to investigate whether our hierarchical category paths perform effectively on documents beyond Wikipedia, we conduct experiment with MS MARCO. Table 2 shows that HyPE consistently improves retrieval accuracy on MS MARCO as well. Although the hierarchical category paths are constructed using Wikipedia category tree as the backbone, the consistent performance gains on MS MARCO emphasize the robustness and generalizability of HyPE. These findings suggest that HyPE can be widely applied to datasets across various domains in the future.
4.3 Hierarchical category paths serve as effective retrieval explanations (RQ2)
We evaluate the explanatory quality of the hierarchical category paths of HyPE through a human evaluation conducted via Amazon Mechanical Turk (AMT). We ask three human judges per sample to compare the quality of the explanations based on four distinct criterias: overall, specificity, reasonability and comprehensiveness. Detailed descriptions of the evaluation criteria and experimental baselines are provided in Appendix A.5.
In Figure 3, HyPE outperforms both the title docid baseline and BM25 across all criteria, receiving high scores for its overall explanation of the retrieval process. Specifically, HyPE shows substantial margin of superiority in terms of specificity and reasonability. This demonstrates that HyPE provides clearer explanations of retrieval process, as well as more logical and reasonable explanation. Furthermore, HyPE beats other baselines in comprehensiveness, indicating that its hierarchical category path is effective in explaining not only narrow, specific details but also broader semantic information. These results highlight that HyPE’s pseudo-reasoning, which utilizes hierarchical category paths, provides users with a effective explanation of the retrieval process.
4.4 HyPE guides users in making better search decision by explanations (RQ3)
In real-world search systems, users are typically provided only with the document title and the first few lines when deciding which result to open. We investigate whether explanations of HyPE can help users effectively identify relevant documents in such real-world settings. To this end, we conduct a human reranking experiment via AMT using the NQ320K dev set. Specifically, human judges rerank the top-5 retrieved results by relevance and rate their confidence (1–5) under two settings: title only, and title with category path. With human-reranking results, we measure performance with Recall@1, MRR@5 and Confidence. Details of the evaluation setup are provided in Appendix A.6.
Table 3 shows that offering hierarchical category paths improve human reranking accuracy, with Recall@1 improving by 23.7% and MRR@5 by 10.4%. This shows that the hierarchical category paths, used as explanations in HyPE, help real-world users better select relevant documents. Additionally, Confidence also improves by 12.0%. These results demonstrate that explanations of HyPE provide users with clarity and guidance, enabling not only more accurate selections but also more confident decisions during search.
Baseline R@1 M@5 Conf.
Title Docid 19.7 47.9 4.0
   + HyPE 24.3 52.8 4.5
Improvement 23.7% 10.4% 12.0%
Table 3: Human reranking performance with and without category paths on NQ320K dev set pairs where the model retrieves the gold document in the top 5.
Document Generated Category Paths for Each Query
Title: Sun The Sun is the star at the center of the Solar System. …The core is the only region of the Sun that produces an appreciable amount of thermal energy through fusion; …The Sun is about halfway through its main-sequence stage, during which nuclear fusion reactions in its core fuse hydrogen into helium. Query 1: the core of the sun in which the sun’s thermonuclear energy is produced takes up about Generated Category Path: universe > energy > energy conversion   Query 2: what stage of the star life cycle is the sun in Generated Category Path: nature > evolution > stellar evolution
Table 4: Example of the document annotated for multiple queries in the NQ320K dev set. The generative retrieval model with HyPE generates query-specific category paths based on the topics of the document associated with each query, explaining why the document is retrieved for the particular query.
5 Analysis
Case Study.
Table 4 illustrates HyPE’s explanations in cases where a single document is annotated with multiple queries on different topics. For the query “the core of the sun in which the sun’s thermonuclear energy is produced”, the model generates paths related to the universe and energy conversion, clearly explaining the thematic relevance between the query and the document. However, for another query, “what stage of the star life cycle is the sun in”, it generates a path related to stellar evolution, which is different from the previously generated path but relevant to the query. This shows that HyPE can provide effective explanations to users by tailoring them to each query.
Figure 4: Performance changes of HyPE. The number of decoded category paths to obtain a ranked docid list.
Analysis of Path-Aware Ranking.
To validate the effectiveness of path-aware ranking strategy, we analyze the performance changes in retrieval accuracy with respect to the number of hierarchical category paths considered by HyPE. Figure 4 presents the analysis results, showing that retrieval accuracy improves as the number of paths increases across all baselines. Notably, there is a clear performance gap between the setting without path-aware ranking strategy (i.e., $K=1$) and with path-aware ranking strategy (i.e., $K>2$). These results indicate that considering multiple paths through the path-aware ranking strategy allows the most relevant docids to be prioritized in the final ranked list, thereby enhancing retrieval accuracy. However, we observe that using too many paths eventually leads to a plateau in performance improvement. Beyond a certain threshold, additional paths tend to introduce noise or increase unnecessary complexity. Consequently, using three paths achieves optimal retrieval accuracy for most docid types.
Analysis of Efficiency.
Providing explanations in the context of generative retrieval inherently increases inference cost, as it involves additional explanation generation beyond the decoding docids alone. Considering this, we conduct additional experiments to analyze the impact of HyPE’s path generation stage on inference cost. Table 5 compares the average inference time per instance for decoding only docids and decoding docids with HyPE’s path generation stage. Details of the analysis setup are provided in Appendix A.8. Overall, when applying HyPE, the inference time increases slightly compared to decoding only docids. Nevertheless, the hierarchical category path employed by HyPE effectively enhances explainability and retrieval accuracy by providing a structured and step-by-step way to convey the connection between queries and retrieved documents, while minimizing the additional computational cost inherently involved in the explanation generation process.
Docid Type Docid Only Docid + HyPE
Summary 0.8127s 0.9134s
Keyword 1.0389s 1.1402s
Table 5: Average inference time per instance for decoding only docid vs decoding both docid and a single path.
6 Related Work
Generative retrieval leverages a single pre-trained generative model, such as T5 Raffel et al. (2020) and BART Lewis et al. (2020), to directly generate document identifier (docid) relevant to the query, enabling end-to-end optimization of the retrieval process Tay et al. (2022); Wang et al. (2022); Sun et al. (2023); Wang et al. (2023); Zhang et al. (2023); Zhou et al. (2022); Lee et al. (2023). Additionally, it reduces reliance on external indexing, lowering the system’s demand for storage resources. However, existing generative retrieval methods directly generate the docid for a user’s query, making it difficult to fully understand why the document is retrieved.
7 Conclusion
In this paper, we propose HyPE, a framework designed to enhance the explainability of document retrieval by utilizing hierarchical category paths. Our experiments demonstrate that HyPE not only enhances overall retrieval performance but also helps users make more accurate decisions during search by providing effective explanations. We hope our research paves the way for meaningful progress in the development of retrieval systems.
Limitations
Despite the promising results and contributions of HyPE, our work has three key limitations stemming from computational costs and budget constraints. First, we do not experiment with alternative backbone hierarchies beyond Wikipedia’s category tree. While it is possible that domain-specific taxonomies may further improve retrieval performance in specialized settings, we consider Wikipedia’s broad and deep hierarchy sufficient for general-purpose document retrieval. Please refer to Appendix A.3 for further discussion. Second, due to cost and scalability constraints, we do not conduct human evaluations to assess how different path depths affect the quality of the explanation. Instead, we provide a limited analysis of explainability with respect to path depth using STS score in the Appendix A.1. Third, we evaluate HyPE using a basic generative retrieval setup (Section 2) to isolate its effect. We do not incorporate advanced optimization techniques or architectures from recent works, which may further improve performance of HyPE.
Ethical Statement
This study strictly adhered to ethical guidelines throughout the human evaluation and data usage process. All content used in the human evaluation and human reranking—including NQ320K and Wikipedia documents—was publicly accessible and did not involve any private or proprietary data. We did not obtain IRB approval for our study, following precedents set by prior work Kim et al. (2023); Kang et al. (2024a) which conducted similar human evaluations without IRB oversight. We ensure that no ethical concerns would arise during the evaluation. The evaluation and reranking were conducted on Amazon Mechanical Turk (AMT), where all participation was anonymous and no personal information was collected at any stage. For human evaluation, we hire three different judges per instance from Amazon Mechanical Turk and guarantee fair compensation for each judge. We pay \$0.15 for each unit task. Human judges were fully informed about the task’s purpose, procedure, and estimated time requirement before beginning the task. Additionally, all examples were screened to exclude offensive, hateful, or sensitive content and were limited to socially and culturally neutral topics. All datasets used in this study are publicly available and appropriately licensed. Specifically, the NQ dataset is distributed under the Apache 2.0 license, and the MS MARCO dataset is released under the MIT license.
Appendix A Appendix
A.1 Quantitative Analysis of Explainability
We quantitatively evaluate whether HyPE’s hierarchical category path provides a valid explanation by effectively capturing the semantic relationship between the query and the document. To this end, we use a semantic textual similarity (STS) model Agirre et al. (2012) [^3] We use sentence-transformers/roberta-base-nli-stsb-mean-tokens as STS model to measure the semantic relevance between two sentences, evaluating the semantic relevance between the query and explanation, as well as between the document and explanation. Specifically, for each baseline, we use the model output as an explanation and calculate the STS scores for both the query-explanation and document-explanation pairs. We then compute the geometric mean of these two scores to evaluate how effectively the explanation captures the relationship between the query and the document. To further analyze the role of hierarchical category paths in explainability, we consider how varying the maximum level of the paths impacts semantic relevance. As mentioned in Section 3.1, HyPE basically leverages Level 4 paths, but we also experiment with varying the maximum level (e.g., Level 2, Level 3) to examine how the maximum level of paths influences the explainability of the query-document relationship. In addition, we also include BM25 as a baseline, which is capable of providing explanations for its retrieval results. For the explanation of BM25, we consider the top-3 terms that have the highest BM25 scores calculated between a given query and a document.
@@ A.1 Quantitative Analysis of Explainability
Table 6: Semantic relevance between query/explanation and document/explanation on 1,000 NQ320K dev set pairs where each baseline successfully retrieves the relevant document at rank 1.